OpenStudy (anonymous):

If g(x) = 3(x − 2), find the value of x if g(x) = 6.

OpenStudy (anonymous):

its 12

OpenStudy (anonymous):

it might help to write it out: g(6) = 3 (6-2) g(6) = 3(4) g(6) = 12